(2t-3)^2+13t=14
Solve for t

Answer by ewatrrr(24785)


Hi
Solving for t
(2t-3)^2 + 13t = 14
4t^2 - 12t + 9 + 13t = 14
4t^2 + t -5 = 0
factoring
(4t + 5)(t-1)= 0 Note: SUM of the inner product(5t) and the outer product(-4t) = t
(4t + 5)=0
t = -5/4
(t-1)= 0
t = 1